OpenStudy (anonymous):

what two integers make y2 = 225 true

OpenStudy (anonymous):

15^2=225

OpenStudy (akashdeepdeb):

a^2 = 225 So a = \(\pm \sqrt{225} = \pm 15\)

OpenStudy (anonymous):

Reason.. Two integers make y2 = 225 You can use This.. y2 = 225 y =15 This is the answer..

OpenStudy (anonymous):

y = -15 y = 15